.(6𝑥2 −𝑦+3)𝑑𝑥+(3𝑦2 −𝑥−2)𝑑𝑦=0


1
Expert's answer
2021-05-24T18:10:22-0400

Ans:-

".(6\ud835\udc65^2 \u2212\ud835\udc66+3)\ud835\udc51\ud835\udc65+(3\ud835\udc66^2 \u2212\ud835\udc65\u22122)\ud835\udc51\ud835\udc66=0"

We consider here the following standard form of ordinary differential equation: "P(x, y)dx + Q(x, y)dy = 0"  

"P(x,y)=6x^2-y+3" and "Q(x,y)=3y^2-x-2"

And we know that "\dfrac{\delta P}{\delta y}=\dfrac{\delta Q}{\delta x}" then

solution of differential equation will be

"\Rightarrow" "du =\dfrac{
∂u}
{∂x} dx +\dfrac
{∂u}
{∂y} dy"

"\Rightarrow P dx + Q dy = 0 ."

One solves "\dfrac{∂u }{∂x} = P" and "\dfrac{∂u}{ ∂y} = Q" to find "u(x, y)"

Now

"\dfrac{∂u}{ ∂y} = 3y^2-x-2\\
u=y^3-(x+2)y+c," "\dfrac{∂u }{∂x} = P\\
u=2x^3-(y-3)x+c"



Then "du = 0" gives "u(x, y) = C," where C is a constant.

"\Rightarrow u(x,y)=2x^3+y^3-(x+2)y-(y-3)x+C"

It will be final solution of the differential equation.